OpenStudy (anonymous):

sharks teeth are shaped like triangles. The height of a tooth of a certain shark is 2 cm longer than the base. the area is 40 cm^2. Find the height and base.

OpenStudy (anonymous):

So area of a triangle is (0.5*base*height)

OpenStudy (anonymous):

Height = Base + 2 cm

OpenStudy (anonymous):

Let base = b so area = (0.5*b*(b+2)) = 40 cm^2

OpenStudy (anonymous):

I'm sorry I just don't get it :/

OpenStudy (anonymous):

okay, so you know that the area of triangle = (1/2)*base*height right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

given that height is 2 cm longer than base right?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So now you can write the area of triangle expression as (1/2)*(base)*(base+2) right? we are replacing height term by base +2 here

OpenStudy (anonymous):

I'm getting lost :(

OpenStudy (anonymous):

okay lets assume a value for the base. How about x cm?

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

so what will be the height?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

nope it will be x+2 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you understood that right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I'm BAD at math :(

OpenStudy (anonymous):

okay no problem, lets start again

OpenStudy (anonymous):

if base = 1 cm what is the height?

OpenStudy (anonymous):

nevermind I give up :/ thanks tho

OpenStudy (anonymous):

umm i can help you out

OpenStudy (anonymous):

if base = 1cm, and height is 2cm longer than base so height should be = 1cm + 2cm right?

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

and similarly if base = 2cm height = 2+2 cm right?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so you see the pattern there?

OpenStudy (anonymous):

a bit

OpenStudy (anonymous):

for every value of base we are just adding 2 to it to get the value of height right?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

you got that right?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

so lets see now you tell me what will be the height if base = 3cm?

OpenStudy (anonymous):

3+3

OpenStudy (anonymous):

or will it be 3+2

OpenStudy (anonymous):

yes it will be 3+2, great!

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

so lets get a step further, what will be the height if the base is x cm?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

No no the base is now x cm long (x is some value you do not know) what will be the height?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

What value will did you add to 3 to get the value of height earlier?

OpenStudy (anonymous):

i have no idea. 2?

OpenStudy (anonymous):

yup 2, thats correct!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so how much will you add to the value of base to get the height?

OpenStudy (anonymous):

You are doing good...

OpenStudy (anonymous):

honestly i dont know

OpenStudy (anonymous):

You were adding 2 right?

OpenStudy (anonymous):

but 2 by what?

OpenStudy (anonymous):

you were adding 2 to the value of base to get the value of height

OpenStudy (anonymous):

uh forget it i dont know this problem. Im frustrated...thank you! bye :)

OpenStudy (anonymous):

okay bye :)